# The Obstacle Problem Arising from the American Chooser Option

### v2

@@ 4.1. Representation of the free boundary
Note that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are closed, and since $V\geqslant J$, we have
$\mathcal{E}_{p}^{\rm ch}\cup\mathcal{E}_{c}^{\rm ch}=\mathcal{E}.$
In the next lemma, we will show that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are disjoint, hence $\partial\mathcal{E}$ is the disjoint union of $\partial\mathcal{E}_{p}^{\rm ch}$ and $\partial\mathcal{E}_{c}^{\rm ch}$. Therefore, examining $\partial\mathcal{E}_{p}^{\rm ch}$ and $\partial\mathcal{E}_{c}^{\rm ch}$ separately is sufficient to analyze $\partial\mathcal{E}$.
We start with proving that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are nonempty.
@@
We first observe from the increasing property of $x_{c}(\tau)$ and the inequality $x_{p}(\tau)<x_{c}(\tau)$ that $C(\tau,x)>e^{x}-K_{c}$ for all $(\tau,x)\in(0,T)\times(-\infty,x_{p}(0))$. Suppose that $\mathcal{E}_{p}^{\rm ch}$ is empty. Choose $x_{0}\in\mathbb{R}$ and $\delta>0$ such that $x_{0}+3\delta<\min\{x_{p}(0),\bar{x}\}$ and set $Q:=(0,\delta)\times(x_{0}-2\delta,x_{0}-\delta)$. Then $V$ satisfies $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$ since
$\partial_{\tau}V-\mathcal{L}V=\begin{cases}\partial_{\tau}C-\mathcal{L}C\quad&\text{in }\ \mathcal{E}\cap Q,\\ 0\quad&\text{in }\ \mathcal{C}\cap Q,\end{cases}$
nd $Q\subset\{C>e^{x}-K_{C}\}$. Therefore,
$(\partial_{\tau}-\mathcal{L})\{V-(K_{p}-e^{x})\}=qe^{x}-rK_{p}\leqslant rK_{p}(e^{-\delta}-1)<0\quad\text{in }\ Q.$
Let $\ell_{Q}:=\{0\}\times(x_{0}-2\delta,x_{0}-\delta)$. By the boundary $C^{\frac{\alpha}{2},\alpha}$ estimate [14, Theorem 6.33] for $\partial^{2}_{x}V$ on $Q\cup\mathcal{P}Q$, we have
$\partial_{\tau}V=\partial_{\tau}V-\mathcal{L}\{V-(K_{p}-e^{x})\}\leqslant rK_{p}(e^{-\delta}-1)<0\quad\text{on }\ \ell_{Q}$
since $V=K_{p}-e^{x}$ on $\ell_{Q}$. This contradicts Lemma 3.5 (1). Thus $\mathcal{E}_{p}^{\rm ch}\neq\emptyset.$ Similarly, we can obtain $\mathcal{E}_{c}^{\rm ch}\neq\emptyset.$
Lemma 4.2.
Let $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ be as in above. Then,
In particular, we have that
and $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are disjoint.
Proof.
Suppose that there exists some point $(\tau_{0},x_{0})\in\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{p}^{\rm ch}$. Since $\{P>K_{p}-e^{x}\}$ is open, there exists some $\delta>0$ such that $Q:=(\tau_{0}-\delta,\tau_{0}]\times(x_{0}-\delta,\infty)\subset\{P>K_{p}-e^{x}\}$ (see Figure 1 (B)). Note that
$\partial_{\tau}(V-P)-\mathcal{L}(V-P)\geqslant 0\quad\text{in }\ Q$
since $\partial_{\tau}P-\mathcal{L}P=0$ in $Q$ and $V-P$ attains its minimum value zero at $(\tau_{0},x_{0})\in Q.$ Thus, by the strong maximum principle [11, Corollary 2.4], $V\equiv P$ in $Q$. This is a contradiction since $C(\tau_{0},x)>P(\tau_{0},x)$ if $x$ is sufficiently large, see Lemma 2.1 (iii), and $V\geqslant C$. Therefore, $\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{p}^{\rm ch}=\emptyset$, i.e., $\mathcal{E}_{p}^{\rm ch}\subset\{P=K_{p}-e^{x}\}$. Using a similar argument, we can obtain $\mathcal{E}_{c}^{\rm ch}\subset\{C=e^{x}-K_{c}\}$.
Since $V\geqslant P\geqslant K_{p}-e^{x}$ and $V\geqslant C\geqslant e^{x}-K_{c}$, it is obvious that $\mathcal{E}_{p}^{\rm ch}\supset\{V=K_{p}-e^{x}\}$ and $\mathcal{E}_{c}^{\rm ch}\supset\{C=e^{x}-K_{c}\}$. This together with (4.1) implies (4.2). Moreover, since $\mathcal{E}_{P}=\{P=K_{p}-e^{x}\}$ and $\mathcal{E}_{C}=\{C=e^{x}-K_{c}\}$ are disjoint (see Subsection 2.4), we conclude that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are also disjoint. ∎
To proceed further, we note that the $x$-coordinate of the points in $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$ is greater than $\ln K_{c}$ and less than $\ln K_{p}$ resepctively (see subsection 2.4). Combined with the above Lemma 4.2, we deduce that the same holds for the $x$-coordinate of each points in $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ respectively. Thus, we take into account Lemma 3.5 and Lemma 4.2 to write the $x$-coordinate of $\partial\mathcal{E}_{p}^{\rm ch}$ and $\partial\mathcal{E}_{c}^{\rm ch}$ by
@@
Remark.
Since $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are nonempty, the set $\{x:V(\tau_{1},x)=K_{p}-e^{x}\}$ and $\{x:V(\tau_{2},x)=e^{x}-K_{c}\}$ are nonempty for some $\tau_{1},\ \tau_{2}\in(0,T)$. Moreover, Lemma 3.5 (1) implies that $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are monotone increasing and decreasing respectively.
@@ 4.2. Properties of the free boundary
At the end of the previous subsection, we represented each free boundary by viewing their $x$-coordinate as the supremum and infimum of the points that touch the obstacle, respectively. To see that such representations correspond to the actual free boundary, we need to establish that the parametrizations in (4.3) and (4.4) are continuous.
Lemma 4.3.
Let $x_{p}^{\rm ch}(\tau),$ $x_{c}^{\rm ch}(\tau):(0,T)\rightarrow\mathbb{R}$ be defined as in (4.3) and (4.4) respectively. Then $x_{p}^{\rm ch}$ and $x_{c}^{\rm ch}$ are continuous on $(0,T)$.
@@
Suppose that $x_{p}^{\rm ch}$ is not continuous at some point $\tau\in(\tau_{0},x_{p}^{\rm ch}(\tau_{0})).$ Then by the monotonicity of $x_{p}^{\rm ch}(\tau),$ either ${\lim_{\tau\to{\tau_{0}}^{+}}x_{p}^{\rm ch}(\tau)}>{x_{p}^{\rm ch}}(\tau_{0})$ or ${\lim_{\tau\to{\tau_{0}}^{-}}x_{p}^{\rm ch}(\tau)}<{x_{p}^{\rm ch}}(\tau_{0})$. Let us consider the first case. Then, there exist some $\varepsilon>0$ and $\delta>0$ such that for all $\tau$ in the interval $(\tau_{0},\tau_{0}+\delta)$ we have
$x_{p}^{\rm ch}(\tau_{0})-x_{p}^{\rm ch}(\tau)\geqslant\varepsilon.$
Note that the cylinder $Q:=(\tau_{0},\tau_{0}+\delta)\times(x_{p}^{\rm ch}(\tau_{0})-\frac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0}))$ is contained in the set $\{V>K_{p}-e^{x}\}$ and not in the set $\mathcal{E}_{c}^{\rm ch}=\{V=C\}$. Hence, $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$ since
$\partial_{\tau}V-\mathcal{L}V=\begin{cases}\partial_{\tau}P-\mathcal{L}P\quad&\text{in}\ \ Q\cap\mathcal{E},\\ 0\quad&\text{in}\ \ Q\cap\mathcal{C},\end{cases}$
and $\partial_{\tau}P-\mathcal{L}P=0$ in $Q\cap\mathcal{E}\subset\{P>K_{p}-e^{x}\}.$ We recall that $x_{p}^{\rm ch}(\tau_{0})<x_{p}^{\rm ch}(0)<\ln\displaystyle{(\frac{r}{q}K_{p})}$ and this yields
@@
Note also that $V=K_{p}-e^{x}$ on the line segment $\ell_{Q}:=\{\tau_{0}\}\times(x_{p}^{\rm ch}(\tau_{0})-\frac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0}))$. Hence, the boundary $C^{\frac{\alpha}{2},\alpha}$ estimate [14, Theorem 6.33] for $\partial^{2}_{x}V$ yields
$\partial_{\tau}V\leqslant qe^{x_{p}^{\rm ch}(\tau_{0})}-rK_{c}+\mathcal{L}\{V-(K_{p}-e^{x})\}<0\quad\text{on}\ \ \ell_{Q}.$
This contradicts Lemma 3.5 (1).
By following the same procedure for the other discontinuity case ${\lim_{\tau\to{\tau_{0}}^{-}}x_{p}^{\rm ch}(\tau)}<{x_{p}^{\rm ch}}(\tau_{0})$, it can be reached at contradiction. Therefore, $x_{p}^{\rm ch}$ is continuous on $(0,T)$. The continuity of $x_{c}^{\rm ch}$ follows analogously. ∎
@@
(i) From Lemma 3.5, we infer that $x_{p}^{\rm ch}(\tau)$ is monotone decreasing. Assume, for contradiction, that $x_{p}^{\rm ch}(\tau)$ is not strictly decreasing. Then there exists some $\tau_{1},\tau_{2}>0$ with $\tau_{1}>\tau_{2}$ such that $x_{p}^{\rm ch}(\tau_{1})=x_{p}^{\rm ch}(\tau_{2})=:x_{0}.$ Note that the cylinder $Q:=(\tau_{1},\tau_{2}-\delta)\times(x_{0},\ln{K_{c}})$ is contained in the continuation region $\mathcal{C}$. This implies $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. Define for each small $\delta>0,$
$V_{\delta}(\tau,x):=V(\tau+\delta,x)-V(\tau,x)\ \ \text{for \ all }\ (\tau,x)\in\Omega_{T-\delta}.$
Then $V_{\delta}$ satisfies
$\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0\quad$ in $\text{in }\ Q,$
$V_{\delta}(\tau,x_{0})=0\quad$ for  all $\text{for \ all }\ \tau\in(\tau_{1},\tau_{2}-\delta).$
@@
Note that there exists a point $(\tau^{*},x^{*})$ such that $(\tau^{*},x^{*})\in Q$ and $V_{\delta}(\tau^{*},x^{*})=0.$ If not, from the parabolic Hopf lemma we infer that
$\partial_{x}V_{\delta}(\tau,x_{0})>0\quad\text{for \ all}\ \ \tau\in(\tau_{1},\tau_{2}-\delta).$
@@
But $V(\tau,x_{0})=K_{p}-e^{x_{0}}$ implies $\partial_{x}V_{\delta}(\tau,x_{0})=0$ for all $\tau\in(\tau_{1},\tau_{2}-\delta)$, which is a contradiction. From Lemma 3.5, $V_{\delta}(\tau,x)\geqslant 0$ for all $(\tau,x)\in\Omega_{T-\delta}$. Thus, $V_{\delta}$ attains its minimum as zero at $(\tau^{*},x^{*})$. Note also that, by the monotonicity of $x_{c}^{\rm ch}(\tau)$ and $x_{p}^{\rm ch}(\tau)$, we can further obtain $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $\mathcal{C}\cap\Omega_{T-\delta}$. Thus, by the Schauder theory it follows that $V_{\delta}$ is smooth in the region $\mathcal{C}\cap\Omega_{T-\delta}$ and applying the strong maximum principle [14, Theorem 2.7] to $V_{\delta}$ yields
(4.6) $V_{\delta}\equiv 0\quad\text{in }\ \mathcal{C}\cap\Omega_{T-\delta}.$
Set a cylinder $Q^{\prime}:=(0,T-\delta)\times(\ln K_{p},\ln K_{c})$. We see (4.3), (4.4) and the Lemma 3.5 to find that the cylinder $Q^{\prime}$ is contained in $\mathcal{C}\cap\Omega_{T-\delta}$. This implies
$V(\tau,x)>\max\{C(\tau,x),P(\tau,x)\}$
for all $(\tau,x)\in Q^{\prime}$. Moreover, from (4.6) and that $\delta>0$ is arbitrary, we deduce that $V$ is constant with respect to $\tau$ in $Q^{\prime}$. Together with $V=\max\{C(0,\cdot),P(0,\cdot)\}$ at the initial boundary of $Q^{\prime}$ and $V\in C(\overline{\Omega_{T}})$, it follows that
$V(\tau,x)=\max\{C(0,x),P(0,x)\}$
for all $(\tau,x)\in Q^{\prime}$. Combining all, we obtain $\max\{C(0,x),P(0,x)\}>\max\{C(\tau,x),P(\tau,x)\}$ for all $(\tau,x)\in Q^{\prime}$, which contradicts $\partial_{\tau}C$ and $\partial_{\tau}P\geqslant 0$ in Lemma 2.1 iii).
(ii) Define $x_{p}^{\rm ch}(0)$ by
$x_{p}^{\rm ch}(0):=\sup\limits\{x:V(0,x)=K_{p}-e^{x}\}.$
Then, since $V(0,x)=\max\{C(0,x),P(0,x)\}$ for all $x\in\mathbb{R}$, it can be verified that $x_{p}^{\rm ch}(0)=\min\{\bar{x},{x_{p}}(0)\}$. We first note that ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}$ exists since $x_{p}^{\rm ch}(\tau)$ is monotone decreasing and bounded above. Suppose ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}>x_{p}^{\rm ch}(0)$. Then there exists some $\tau_{p}>0$ such that $x_{p}^{\rm ch}(\tau_{p})>x_{p}^{\rm ch}(0)$. This implies
$K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})}$ $=V(\tau_{p},x_{p}^{\rm ch}(\tau_{p}))\geqslant J(\tau_{p},x_{p}^{\rm ch}(\tau_{p}))$
$\geqslant J(0,x_{p}^{\rm ch}(\tau_{p}))>K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})},$
@@
where the second inequality is due to $\partial_{\tau}C\geqslant 0$ and $\partial_{\tau}P\geqslant 0$ (see Lemma 2.1 iii)). This is a contradiction.
@@
On the other hand, if ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}$ is strictly less than $x_{p}^{\rm ch}(0)$, there exist some $\epsilon>0$ and $\delta>0$ such that for all $\tau\in(0,\delta)$,
@@
Now, consider a cylinder $Q:=(0,\delta)\times(x^{\rm ch}_{p}(0)-\frac{\varepsilon}{2},x^{\rm ch}_{p}(0))$. Since $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$ and $x^{\rm ch}_{p}(0)<\ln(\displaystyle{\frac{r}{q}}K_{p})$,

### v3

@@ 4.1. Representation of the free boundary
In the following lemma, we prove that $V$ contacts the obstacle exclusively along its smooth portions; that is, $V$ coincides with either $e^{x}-K_{c}$ or $K_{p}-e^{x}$ in the exercise region $\mathcal{E}$. This ensures that standard arguments remain applicable within our framework.
Furthermore, we establish that $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are mutually disjoint. Consequently, the free boundary $\partial\mathcal{E}$ decomposes into the disjoint union of $\partial\mathcal{E}_{P}^{\rm ch}$ and $\partial\mathcal{E}_{C}^{\rm ch}$. Therefore, it suffices to analyze $\partial\mathcal{E}_{P}^{\rm ch}$ and $\partial\mathcal{E}_{C}^{\rm ch}$ separately.
Lemma 4.1.
Let $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ be defined as above. Then,
In particular, we have
which further implies that $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are mutually disjoint.
Proof.
Suppose that there exists a point $(\tau_{0},x_{0})\in\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{P}^{\rm ch}$. Since $\{P>K_{p}-e^{x}\}$ is open, there exists some $\delta>0$ such that
$Q:=(\tau_{0}-\delta,\tau_{0}]\times(x_{0}-\delta,\infty)\subset\{P>K_{p}-e^{x}\}$
(see Figure 1 (B)). Then, by the above inclusion, it follows that $\partial_{\tau}P-\mathcal{L}P=0$ in $Q$ and thus,
$\partial_{\tau}(V-P)-\mathcal{L}(V-P)\geqslant 0\quad\text{a.e. in }\ Q.$
Moreover, since $(\tau_{0},x_{0})\in\mathcal{E}_{P}^{\rm ch}$, VPV-P attains its minimum value 0 at the interior point $(\tau_{0},x_{0})\in Q$. Thus, by the strong maximum principle [18, Corollary 2.4], we deduce that $V\equiv P$ in $Q$. This is a contradiction since $C(\tau_{0},x)>P(\tau_{0},x)$ for all $x>x_{\tau_{0}}$ (see Lemma 2.1 (iv)) and $V\geqslant C$ in $\Omega_{T}$. Therefore,
$\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{P}^{\rm ch}=\emptyset.$
In other words, $\mathcal{E}_{P}^{\rm ch}$ is contained in the set $\mathcal{E}_{P}=\{P=K_{p}-e^{x}\}$. Using a similar argument with $V$ and $C$, we can also obtain $\mathcal{E}_{C}^{\rm ch}\subset\mathcal{E}_{C}$.
Since $V\geqslant P\geqslant K_{p}-e^{x}$ and $V\geqslant C\geqslant e^{x}-K_{c}$, we have
$\{V=K_{p}-e^{x}\}\subset\mathcal{E}_{P}^{\rm ch}\quad\text{and}\quad\{V=e^{x}-K_{c}\}\subset\mathcal{E}_{C}^{\rm ch}.$
Combined with (4.1), these inclusions yield (4.2). Furthermore, since $\mathcal{E}_{P}$ and $\mathcal{E}_{C}$ are disjoint (see Subsection 2.4), the inclusions $\mathcal{E}_{P}^{\rm ch}\subset\mathcal{E}_{P}$ and $\mathcal{E}_{C}^{\rm ch}\subset\mathcal{E}_{C}$ immediately imply the disjointness of $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$. ∎
@@
We first observe from the increasing property of $x_{c}(\tau)$ and the inequality $x_{p}(\tau)<x_{c}(\tau)$ that $C(\tau,x)>e^{x}-K_{c}$ for all $(\tau,x)\in(0,T)\times(-\infty,x_{p}(0))$. Suppose, by contradiction, that $\mathcal{E}_{P}^{\rm ch}=\emptyset$. Choose $x_{0}\in\mathbb{R}$ and $\delta>0$ such that $x_{0}+3\delta<\min\{x_{p}(0),\bar{x}\}$, and set
$Q:=(0,\delta)\times(x_{0}-2\delta,x_{0}-\delta).$
Since $Q\subset\{C>e^{x}-K_{c}\}$ and the set $\{V=C\}$ is identical to $\{V=e^{x}-K_{c}\}$ by Lemma 4.1, it follows that $Q$ is disjoint from $\{V=C\}$. Furthermore, under the assumption that $\mathcal{E}_{P}^{\rm ch}=\emptyset$, the set $Q$ is necessarily contained within the continuation region $\mathcal{C}$. Consequently, we obtain $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$.
Define the bottom boundary of $Q$ as
$\mathcal{B}_{Q}:=\{0\}\times(x_{0}-2\delta,x_{0}-\delta).$
Then from the initial condition $V(0,x)=\max\{C(0,x),P(0,x)\}$ and $x_{0}<\min\{x_{p}(0),\bar{x}\}$, it follows that $V(0,x)=K_{p}-e^{x}$ on $\mathcal{B}_{Q}$. Moreover, since $V$ satisfies the equation $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$, we apply the result in [8, Theorem 19, p. 321] to $V-(K_{p}-e^{x})$ to deduce that $V$ is smooth in $Q$ and up to the initial boundary $\mathcal{B}_{Q}$. Therefore, passing to the limit as $\tau\to 0^{+}$ in $(\partial_{\tau}-\mathcal{L})(V-(K_{p}-e^{x}))=qe^{x}-rK_{p}$, we obtain for sufficiently small $\delta>0$ that
$\partial_{\tau}V\leqslant rK_{p}(e^{-\delta}-1)<0\quad\text{on }\mathcal{B}_{Q}.$
This contradicts the first estimate in Lemma 3.5. Hence, $\mathcal{E}_{P}^{\rm ch}\neq\emptyset$.
The same argument yields $\mathcal{E}_{C}^{\rm ch}\neq\emptyset$. ∎
To proceed further, recall from Subsection 2.4 that the $x$-coordinates of points in $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$ satisfy $x>\ln K_{c}$ and $x<\ln K_{p}$, respectively. Combined with Lemma 4.1, the same bounds hold for points in $\mathcal{E}_{C}^{\rm ch}$ and $\mathcal{E}_{P}^{\rm ch}$, respectively. Therefore, using Lemma 3.5 and Lemma 4.1, we parameterize the free boundaries by
@@
Remark 4.3.
Since $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are nonempty, the set $\{x:V(\tau_{1},x)=K_{p}-e^{x}\}$ and $\{x:V(\tau_{2},x)=e^{x}-K_{c}\}$ are nonempty for some $\tau_{1},\ \tau_{2}\in(0,T)$. Moreover, the monotonicity of $V$ with respect to $\tau$, obtained in Lemma 3.5, implies that $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are monotone decreasing and increasing respectively.
@@ 4.2. Properties of the free boundary
At the end of the previous subsection, we represented each free boundary by viewing their $x$-coordinate as the supremum and infimum of the points where the contact between the solution and the obstacle occurs. To see that such representations correspond to the actual free boundary, we need to establish that the parametrizations in (4.3) and (4.4) are continuous.
Lemma $4.$
Let $x_{p}^{\rm ch}(\tau),$ $x_{c}^{\rm ch}(\tau)$ be defined as in (4.3) and (4.4) respectively. Then $x_{p}^{\rm ch}$ and $x_{c}^{\rm ch}$ are continuous on $(0,T)$.
@@
Suppose that $x_{p}^{\rm ch}$ is not continuous at some point $\tau_{0}\in(0,T).$ Then by the monotonicity of $x_{p}^{\rm ch}(\tau),$ either ${\lim_{\tau\to{\tau_{0}}^{+}}x_{p}^{\rm ch}(\tau)}>{x_{p}^{\rm ch}}(\tau_{0})$ or ${\lim_{\tau\to{\tau_{0}}^{-}}x_{p}^{\rm ch}(\tau)}<{x_{p}^{\rm ch}}(\tau_{0})$. Let us consider the first case. Then, there exist some $\varepsilon>0$ and $\delta>0$ such that for all $\tau$ in the interval $(\tau_{0},\tau_{0}+\delta)$ we have
$x_{p}^{\rm ch}(\tau)-x_{p}^{\rm ch}(\tau_{0})\geqslant\varepsilon.$
Consider the cylinder
$Q:=(\tau_{0},\tau_{0}+\delta)\times\big(x_{p}^{\rm ch}(\tau_{0})-\tfrac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0})\big).$
By construction, it can be seen that $Q$ is contained in the set $\{V>K_{p}-e^{x}\}$ and disjoint from $\{V=e^{x}-K_{c}\}$. Hence, $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. Moreover, since $V=K_{p}-e^{x}$ on the bottom boundary $\mathcal{B}_{Q}:=\{\tau_{0}\}\times(x_{p}^{\rm ch}(\tau_{0})-\frac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0}))$ of $Q$, we apply the result in [8, Theorem 19, p. 321] to $V-(K_{p}-e^{x})$ to deduce that $V$ is smooth in $Q$ up to the initial boundary $\mathcal{B}_{Q}$. Next, recall from the monotonicity of $x_{p}^{\rm ch}$ that $x_{p}^{\rm ch}(\tau_{0})<x_{p}^{\rm ch}(0)\leqslant\ln{(\frac{r}{q}K_{p})}$ and this implies
@@
Using the previous regularity result, letting $\tau\to\tau_{0}^{+}$ yields
$\partial_{\tau}V\leqslant qe^{x_{p}^{\rm ch}(\tau_{0})}-rK_{p}<0\quad\text{on }\ \mathcal{B}_{Q}.$
This contradicts the first estimate in Lemma 3.5.
The same argument applies to the other discontinuity case
$\lim_{\tau\to\tau_{0}^{-}}x_{p}^{\rm ch}(\tau)<x_{p}^{\rm ch}(\tau_{0}),$
which also leads to a contradiction. Therefore, $x_{p}^{\rm ch}$ is continuous on $(0,T)$. The continuity of $x_{c}^{\rm ch}$ follows analogously. ∎
@@
From Lemma 3.5, we can easily verify that $x_{p}^{\rm ch}(\tau)$ is monotone decreasing. Assume, for contradiction, that $x_{p}^{\rm ch}(\tau)$ is not strictly decreasing. Then there exists some $\tau_{1},\tau_{2}>0$ with $\tau_{1}<\tau_{2}$ such that $x_{p}^{\rm ch}(\tau_{1})=x_{p}^{\rm ch}(\tau_{2})=:x_{0}.$ Moreover, from that $x_{c}^{\rm ch}(0)\geqslant x_{c}(0)>\ln K_{c}$ (see subsection 2.4 and (4.5)), we see the cylinder $Q:=(\tau_{1},\tau_{2}-\delta)\times(x_{0},\ln{K_{c}})$ is contained in the continuation region $\mathcal{C}$. This implies $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. For each sufficiently small $\delta>0$, we define by
$V_{\delta}(\tau,x):=V(\tau+\delta,x)-V(\tau,x)\quad\text{for all }\ (\tau,x)\in\Omega_{T-\delta}.$
@@
Then, $V_{\delta}$ satisfies $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $Q$. Note that there exists a point $(\tau^{\prime},x^{\prime})\in Q$ such that $V_{\delta}(\tau^{\prime},x^{\prime})=0.$ Otherwise, we see from the parabolic Hopf lemma that
(4.6) $\partial_{x}V_{\delta}(\tau,x_{0})>0\quad\text{for \ all}\ \ \tau\in(\tau_{1},\tau_{2}-\delta).$
But $V(\tau,x_{0})=K_{p}-e^{x_{0}}$ and the continuity of $\partial_{x}V$, induced by the Sobolev embedding theorem, implies that $\partial_{x}V_{\delta}(\tau,x_{0})=0$ for all $\tau\in(\tau_{1},\tau_{2}-\delta)$, which contradicts (4.6).
@@
Now, by the estimate in Lemma 3.5 and the continuity of $V$, it follows that $V_{\delta}(\tau,x)\geqslant 0$ for all $(\tau,x)\in Q$. Thus, $V_{\delta}$ attains its minimum as zero at $(\tau^{\prime},x^{\prime})\in Q$. Furthermore, since $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $Q$, we deduce from the Schauder theory that $V_{\delta}$ is smooth in $Q$. Applying the strong maximum principle [17, Theorem 2.7] to $V_{\delta}$ yields
(4.7) $V_{\delta}=0\quad\text{in }\ \{(\tau,x)\in Q:\tau\leqslant\tau^{\prime}\}.$
Define the cylinder $Q^{\prime}:=(0,\tau^{\prime})\times(\ln K_{p},\ln K_{c})$. In view of (4.3), (4.4), and the monotonicity of each free boundaries obtained by the estimates in Lemma 3.5, we find that $Q^{\prime}\subset\mathcal{C}\cap\Omega_{T-\delta}$, and this implies $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $Q^{\prime}$. Moreover, since (4.7) ensures that $V_{\delta}\equiv 0$ in the non-empty set $Q\cap Q^{\prime}$, applying the strong maximum principle again yields $V_{\delta}\equiv 0$ in $Q^{\prime}$. Since $\delta>0$ is arbitrary, we conclude that $V$ is constant with respect to $\tau$ in $Q^{\prime}$. Considering the initial condition $V(0,x)=\max\{C(0,x),P(0,x)\}$ and the regularity $V\in C(\overline{\Omega_{T}})$, it follows that
(4.8) $V(\tau,x)=\max\{C(0,x),P(0,x)\}\quad\text{for all }(\tau,x)\in Q^{\prime}.$
Since $C$ and $P$ are non-decreasing with respect to $\tau$ by Lemma 2.1 (iii), we have
$\max\{C(0,x),P(0,x)\}\leqslant\max\{C(\tau,x),P(\tau,x)\}\quad\text{for all }(\tau,x)\in Q^{\prime}.$
However, the inclusion $Q^{\prime}\subset\mathcal{C}$ yields
$V(\tau,x)>\max\{C(\tau,x),P(\tau,x)\}\geqslant\max\{C(0,x),P(0,x)\}\quad\text{for all }(\tau,x)\in Q^{\prime},$
and this contradicts (4.8) and completes the proof. ∎
Proof of ii).
We define $x_{p}^{\rm ch}(0)$ by
$x_{p}^{\rm ch}(0):=\sup\{x\in\mathbb{R}:V(0,x)=K_{p}-e^{x}\}.$
Since $V(0,x)=\max\{C(0,x),P(0,x)\}$ for all $x\in\mathbb{R}$, it is straightforward to verify that
$x_{p}^{\rm ch}(0)=\min\{\bar{x},x_{p}(0)\}.$
We first observe that $\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)$ exists, since $x_{p}^{\rm ch}(\tau)$ is monotonically decreasing and bounded from above.
Suppose ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}>x_{p}^{\rm ch}(0)$. Then there exists some $\tau_{p}>0$ such that $x_{p}^{\rm ch}(\tau_{p})>x_{p}^{\rm ch}(0)$. Hence,
$K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})}=V\big(\tau_{p},x_{p}^{\rm ch}(\tau_{p})\big)\geqslant J\big(\tau_{p},x_{p}^{\rm ch}(\tau_{p})\big)\geqslant J\big(0,x_{p}^{\rm ch}(\tau_{p})\big)>K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})},$
@@
where $J=\max\{C,P\}$ and the second inequality follows from the result $\partial_{\tau}C\geqslant 0$ and $\partial_{\tau}P\geqslant 0$ in $\Omega_{T}$ (see Lemma 2.1 iii)). Therefore, a contradiction occurs.
@@
On the other hand, if ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}$ is strictly less than $x_{p}^{\rm ch}(0)$, there exist some $\varepsilon>0$ and $\delta>0$ such that for all $\tau\in(0,\delta)$,
@@
Now, consider a cylinder $Q:=(0,\delta)\times(x^{\rm ch}_{p}(0)-\frac{\varepsilon}{2},x^{\rm ch}_{p}(0))$. By construction, we see that $Q$ is contained in the continuation region $\mathcal{C}$ and thus, $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. Moreover, since $x^{\rm ch}_{p}(0)<\ln({\frac{r}{q}}K_{p})$, it follows that